Review If the electric charge on the object is doubled while its mass remains the same, find the direction and magnitude of its acceleration. Express your answer using three significant figures. An object with a charge of -3.8 AC and a mass of 4.7x10-2 kg experiences an upward electric force, due to a uniform electric field, equal in magnitude to its weight. Submit Request Answer Part D If the electric charge on the object is doubled while its mass remains the same, find the direction and magnitude of its acceleration. upward downward to the left to the right O Submit Request Answer
Answer #1

Ans:-

Part C.

m=4.7x10^-2 Kg.

Electric charge of the object(q)=2*3.8μC

=7.6μC

Acceleration(a)=Eq/m

Electric Field(E)=mg/q

so,

a=(mg/q)*(q/m)

=g

=9.81m/s2 (in the direction of force, i,e-upward)

hence,

a=9.81 m/sec2 ...ans

Part D.

ans:- Upward.